# Bridging Discrete and Continuous: A Multimodal Strategy for Complex Emotion Detection

### v1

@@ Abstract
Index Terms:
Multimodal emotion recognition, Emotional variability, Valence-Arousal-Dominance (VAD) framework, Machine learning in emotion detection
I Introduction
Human emotions are complex and described through diverse vocabularies across languages, reflecting our thoughts, feelings, and reactions via facial expressions, body language, voice tone, and speech. [1]. Accurate comprehension and response to human emotions by machines can significantly benefit areas such as marketing, music, mental health monitoring, and human-computer interaction [2, 3, 4, 5]. Thus, developing systems capable of precisely identifying varieties of human emotions is essential.
@@ I Introduction
Figure 1: Emotion Vocabularies in 3D VAD Space
@@
The experimental results demonstrate that the framework successfully achieved the transformation between discrete and continuous models. By comparing their evaluation metrics, we conclude that the proposed framework performs the transformation with a comparable level of accuracy. We also compared the similarity score between the generated open emotion vocabulary set and the original dataset’s open set, achieving a high score that indicates significant semantic overlap. This shows that our model can reliably produce a nuanced and flexible emotion vocabulary, which could be effectively applied to tasks requiring more detailed emotion recognition and analysis.
The contribution of this paper is summarized as follows:
@@
II Related work
Multimodal emotion detection methods primarily differ in two key areas: the approaches to categorizing and measuring emotions, and the techniques used to extract and integrate features from various modalities.
II-A The measurement of emotion
@@
II-B Multimodal Emotion Detection
@@
Visual, audio, and textual features are captured separately using different methods. Visual features primarily focus on facial expressions, which can be seen as facial muscle movements. The Facial Action Coding System (FACS) [20] has been developed to manually code facial expressions using action units (AUs) based on specific muscle movements. Inspired by their work, many researchers have utilized image and video processing to analyze and categorize facial expressions by tracking features and measuring movement, applying these methods to the “basic expressions” identified in multimodal emotion studies [21, 22, 23, 24]. Similar to other audio domains [25, 26, 27], emotion-related features consist of two main categories: linguistic information, which pertains to the content of speech, and paralinguistic information, which captures emotions through the tone and manner of delivery [28]. In addition to commonly used models like CNN, RNN, and LSTM [29, 30], other approaches also make promising prediction. For instance, the use of features such as mel-frequency cepstral coefficient (MFCC), perceptual linear prediction coefficient (PLPC), and perceptual minimum variance distortionless response (PMVDR) coefficient [31] . Textual emotions are expressed through speech and can be abstract by transcripts. Traditional techniques such as the bag-of-words (BoW) model [32, 33] have been commonly used. Then word embedding methods like tf-idf, word2vec [34] and GloVe [35] were based mainly on syntactic contexts. For handling larger text, BERT (Bidirectional Encoder Representations from Transformers) uses transformer-based encoders that assess both preceding and subsequent contexts to better predict words in sentences [36].
III Methodology
The project aims to bridge discrete and continuous emotion systems, using a multimodal model to detect emotions and produce more nuanced emotion labels beyond the basic six. This approach allows for working with diverse datasets and improves the variety and depth of emotion detection.
@@ III Methodology
We first transformed discrete emotion labels into continuous emotion labels, then applied and refined several models for the task. The first model, Multimodal End-to-End Sparse Model (ME2E) [37], is a multimodal emotion detection model that uses discrete emotion labels as input and outputs corresponding discrete emotions, serving as our baseline. The second, ME2E Lite model, a refined version of ME2E to better align with our dataset for improved performance. The third, Proposed VAD, builds on the ME2E Lite pipeline but uses 3D VAD score as input, outputting continuous VAD scores that can be mapped to discrete emotions while also generating a broader range of nuanced emotion vocabularies.
@@
III-A The K-means clustering classifier - Transformation between discrete emotion labels and continuous emotion labels
To transition from discrete to multidimensional emotion methods, we developed a classifier that enables the transformation between discrete and continuous emotion labels using the NRC-VAD lexicon [38].
@@
We extracted 195 emotion vocabularies with assigned VAD scores from the 20,000-vocabularies NRC-VAD lexicon, using the polar term based on a -1 to 1 scale, as taking the absolute values of these scores is likely to yield better features than those derived from a 0 to 1 scale. The six basic emotions selected—happy, sad, worried, surprised, angry, and neutral—align with the categories used in our dataset. As VAD scores are available for all selected basic emotions except neutral, we assigned a VAD score of (0,0,0) to neutral, reflecting a lack of significant emotional fluctuations typically associated with this state.
@@
This approach effectively allows us to bridge continuous and discrete emotional labels, with the results of the clustering visualized in Figure 1 and Table I, demonstrating how similar emotions are grouped together based on their VAD scores.
@@
TABLE I: Emotion Clustering
Emotion Name Sampled Emotion in Cluster
@@ III-B ME2E - Baseline Model
III-C ME2E Lite - The refined Model
@@
III-D Proposed VAD Model – The model bridges the gap between discrete and continuous emotions
@@
IV Dataset
We have selected the MER2024 dataset [43] for our study. This dataset consists of raw video clips collected from Chinese movies and TV series, ensuring cultural consistency. The raw video samples were split into smaller video clips, each containing a complete segment from the same character. The dataset has been labeled with six discrete emotions: happy, angry, sad, neutral, worried, and surprised. Additionally, there is a subset of open-vocabulary labels and explanations [44], which allows us to evaluate a more dynamic emotion output. Each sample contains audio data with a sampling rate of 44.1 kHz, text transcript and video frames. The video operates at 25 frames per second (FPS), and the frames are sampled every 500 milliseconds which result in capturing about two frames each second.We created a dataloader to split the data into training (70%), validation (15%), and test sets (15%).The details of our data split are shown in Table II.
TABLE II: Dataset Distribution
Emotion Total Train Val Test Clip Avg Words Avg
@@ IV Dataset
V Experiments setup and Evaluation
@@
The Proposed VAD model We used a batch size of 8, reaching optimal performance with a learning rate of 0.0005 and a weight decay of 0.0001. Since it is a continuous task, we used Mean Squared Error (MSE) as the loss function and evaluation metric.We also evaluate using L2 distance, Mean Absolute Error (MAE), and Pearson Correlation Coefficient (PCC) to assess the linear relationship between predictions and actual values, where $\text{PCC}=\frac{\sum_{i=1}^{N}(y_{i}-\mu_{y})(\hat{y}_{i}-\mu_{\hat{y}})}{\sqrt{\sum_{i=1}^{M}(y_{i}-\mu_{y})^{2}}\cdot\sqrt{\sum_{i=1}^{M}(\hat{y}_{i}-\mu_{\hat{y}})^{2}}}$.
VI Results Analysis
@@
VI-A Model performance on continuous emotion detection
Table III illustrates the performance of proposed VAD model which deals with the continuous VAD emotion labels. The L2 distance of 0.64 suggests that the model’s predictions are relatively close to the actual values, while the MSE of 0.19 and MAE of 0.36 indicate that the model’s predictions are generally accurate. The PCC of 0.47 further confirms that the model has learned to predict the emotional values and has achieved a reasonable level of correlation between the predicted and actual values.
TABLE III: Result on Continuous Emotion Detection
@@ VI-A Model performance on continuous emotion detection
VI-B Model performance comparison on discrete emotion detection
@@
Table IV shows that the the proposed VAD model stands out with higher precision (0.49) and recall (0.45), suggesting it is more effective in accurately identifying and capturing emotions. ME2E Lite model also performs well, with an F1 score of 0.42, and balanced precision and recall at 0.42 and 0.43, respectively. The baseline ME2E model had the lowest performance, with an F1 score of 0.33 and lower precision and recall (both at 0.32).
@@
TABLE IV: Result on Discrete Emotion Detection
@@
VI-C Open-vocabulary exploration
At the final stage of our study, we explored the possibility of generating open-vocabulary emotional responses based on six discrete emotion classes. For this, we utilized a subset of 68 samples from the MER2024 dataset which contains the open-vocabulary outputs. We then applied our proposed VAD model to this subset and mapped the resulting VAD labels into the 3D emotion word space showed in the Introduction.
To output the final dataset, we used an L2 distance threshold of 0.25, calculated based on the distance which could output an average of five emotion vocabularies from our 3D Emotion Vocabulary Space.
Our finds (see supplementary material for details) indicate that our model effectively predicts nuanced emotional states. For instance, in Sample 00000368, the MER2024 dataset listed emotions like ’Alert,’ ’Excited,’ ’Confused,’ and ’Curious,’ and our model predicted ’Shocked.’ This aligns well as ’Shocked’ can encapsulate alertness, excitement, and confusion. In Sample 0002419, the dataset included ’Calm,’ ’Relaxed,’ and ’Happy,’ while our model suggested ’Caring’ and ’Curious.’ These predictions are compatible, with ’Caring’ reflecting a relaxed and content state. Full output list of To further examine the alignment of our open vocabulary, we assess the semantic similarity across various model’s output of emotion labels.
Similarity between open-set vocabularies: We employed the pre-trained Word2Vec model from Google News [45]. This model enabled us to compare the word sets by converting them into vector representations and calculating their cosine similarity based on of the vector space embeddings.
Figure 3: Emotion Vocabularies in 3D VAD Space
We began by assessing the similarity between EMER24’s discrete emotion ground truth and the EMER24 OPEN emotion set, achieving a mean similarity of approximately 0.879. This establishes a benchmark, indicating that ideal model performance should aim for this level to ensure consistent ground truth alignment.
We then evaluated the similarity between the EMER24 OPEN set and the proposed VAD open set, resulting in a lower mean value of 0.85. This decrease was mainly due to outliers with a similarity of 0, stemming from an insufficient emotion vocabulary in the 3D emotion space affecting the VAD score distribution, due to the distribution of emotion lexicon. Despite this, the median similarity in this comparison is higher, illustrating a strong central tendency that supports the VAD set’s reliability.
Lastly, we compared the EMER24 OPEN set with the ME2E Lite model’s predictions, observing a slightly higher mean similarity of 0.86. This suggests a marginally better alignment compared to the VAD set, indicating that the ME2E Lite model also produces a consistent and reliable outputs.
VII Conclusion

### v2

@@ Abstract
Index Terms—  Multimodal emotion recognition, emotional variability, valence-arousal-dominance (VAD) framework, emotion detection, machine learning
1 Introduction
Human emotions are complex and described through diverse vocabularies across languages, reflecting our thoughts, feelings, and reactions via facial expressions, body language, voice tone, and speech. [1]. Accurate comprehension and response to human emotions by machines can significantly benefit areas such as marketing, mental health monitoring, multimedia generation, and human-computer interaction [2, 3, 4, 5, 6, 7]. Thus, developing systems capable of precisely identifying varieties of human emotions is essential.
@@ 1 Introduction
Fig. 1: Emotion Vocabularies in 3D VAD Space
@@
The experimental results demonstrate that the framework successfully achieved the transformation between discrete and continuous models. By comparing their evaluation metrics, we conclude that the proposed framework performs the transformation with a comparable level of accuracy. We also compared the similarity score between the generated open emotion vocabulary set and the original dataset’s open set, achieving a high score that indicates significant semantic overlap. This shows that our model can reliably produce a nuanced and flexible emotion vocabulary, which could be effectively applied to tasks requiring more detailed emotion recognition and analysis. The contribution of this paper is summarized as follows:
@@
2 Related work
Compared to regular classification tasks [10, 11, 12], multimodal emotion detection methods primarily differ in two key areas: the approaches to categorizing and measuring emotions, and the techniques used to extract and integrate features from various modalities.
2.1 The measurement of emotion
@@
2.2 Multimodal Emotion Detection
@@
Visual, audio, and textual features are captured separately using different methods. Visual features primarily focus on facial expressions, which can be seen as facial muscle movements. The Facial Action Coding System (FACS) [25] has been developed to manually code facial expressions using action units (AUs) based on specific muscle movements. Inspired by their work, many researchers have utilized image and video processing to analyze and categorize facial expressions by tracking features and measuring movement, applying these methods to the “basic expressions” identified in multimodal emotion studies [26, 27, 28, 29]. Similar to other audio domains [30, 7], emotion-related features consist of two main categories: linguistic information, which pertains to the content of speech, and paralinguistic information, which captures emotions through the tone and manner of delivery [31]. In addition to commonly used models like CNN, RNN, and LSTM [32, 33], other approaches also make promising prediction. For instance, the use of features such as mel-frequency cepstral coefficient (MFCC), perceptual linear prediction coefficient (PLPC), and perceptual minimum variance distortionless response (PMVDR) coefficient [34] . Textual emotions are expressed through speech and can be abstract by transcripts. Traditional techniques such as the bag-of-words (BoW) model [35, 36] have been commonly used. Then word embedding methods like tf-idf, word2vec [37] and GloVe [38] were based mainly on syntactic contexts. For handling larger text, BERT (Bidirectional Encoder Representations from Transformers) uses transformer-based encoders that assess both preceding and subsequent contexts to better predict words in sentences [39].
3 Methodology
@@
The project aims to bridge discrete and continuous emotion systems, using a multimodal model to detect emotions and produce more nuanced emotion labels beyond the basic six. This approach allows for working with diverse datasets and improves the variety and depth of emotion detection. We first transformed discrete emotion labels into continuous emotion labels, then applied and refined several models for the task. The first model, Multimodal End-to-End Sparse Model (ME2E) [40], is a multimodal emotion detection model that uses discrete emotion labels as input and outputs corresponding discrete emotions, serving as our baseline. The second, ME2E Lite model, a refined version of ME2E to better align with our dataset for improved performance. The third, Proposed VAD, builds on the ME2E Lite pipeline but uses 3D VAD score as input, outputting continuous VAD scores that can be mapped to discrete emotions while also generating a broader range of nuanced emotion vocabularies.
@@
3.1 The K-means clustering classifier - Transformation between discrete emotion labels and continuous emotion labels
@@
To transition from discrete to multidimensional emotion methods, we developed a classifier that enables the transformation between discrete and continuous emotion labels using the NRC-VAD lexicon [41]. We extracted 195 emotion vocabularies with assigned VAD scores from the 20,000-vocabularies NRC-VAD lexicon, using the polar term based on a -1 to 1 scale, as taking the absolute values of these scores is likely to yield better features than those derived from a 0 to 1 scale. The six basic emotions selected—happy, sad, worried, surprised, angry, and neutral—align with the categories used in our dataset. As VAD scores are available for all selected basic emotions except neutral, we assigned a VAD score of (0,0,0) to neutral, reflecting a lack of significant emotional fluctuations typically associated with this state.
@@
This approach effectively allows us to bridge continuous and discrete emotional labels, with the results of the clustering visualized in Figure 1 and Table 1, demonstrating how similar emotions are grouped together based on their VAD scores.
@@
Table 1: Emotion Clustering
Cluster Examples in Cluster
@@ 3.2 ME2E - Baseline Model
3.3 ME2E Lite - The refined Model
@@
3.4 Proposed VAD Model – The model bridges the gap between discrete and continuous emotions
@@
4 Dataset
We have selected the MER2024 dataset [46] for our study. This dataset consists of raw video clips collected from Chinese movies and TV series, ensuring cultural consistency. The raw video samples were split into smaller video clips, each containing a complete segment from the same character. The dataset has been labeled with six discrete emotions: happy, angry, sad, neutral, worried, and surprised. Additionally, there is a subset of open-vocabulary labels and explanations [47], which allows us to evaluate a more dynamic emotion output. Each sample contains audio data with a sampling rate of 44.1 kHz, text transcript and video frames. The video operates at 25 frames per second (FPS), and the frames are sampled every 500 milliseconds which result in capturing about two frames each second.We created a dataloader to split the data into training (70%), validation (15%), and test sets (15%).The details of our data split are shown in Table 2.
Table 2: Dataset Distribution. C-Avg and W-Avg represent the average of clip duration and word counts, respectively.
Emotion Total Train Val Test C-Avg W-Avg
@@ 4 Dataset
5 Experiments setup and Evaluation
@@
The Proposed VAD model We used a batch size of 8, reaching optimal performance with a learning rate of 0.0005 and a weight decay of 0.0001. Since it is a continuous task, we used Mean Squared Error (MSE) as the loss function and evaluation metric.We also evaluate using L2 distance, Mean Absolute Error (MAE), and Pearson Correlation Coefficient (PCC) to assess the linear relationship between predictions and actual values, where
$\text{PCC}=\frac{\sum_{i=1}^{N}(y_{i}-\mu_{y})(\hat{y}_{i}-\mu_{\hat{y}})}{\sqrt{\sum_{i=1}^{M}(y_{i}-\mu_{y})^{2}}\cdot\sqrt{\sum_{i=1}^{M}(\hat{y}_{i}-\mu_{\hat{y}})^{2}}}.$
6 Results Analysis
@@
6.1 Model performance on continuous emotion detection
Table 3 illustrates the performance of proposed VAD model which deals with the continuous VAD emotion labels. The L2 distance of 0.64 suggests that the model’s predictions are relatively close to the actual values, while the MSE of 0.19 and MAE of 0.36 indicate that the model’s predictions are generally accurate. The PCC of 0.47 further confirms that the model has learned to predict the emotional values and has achieved a reasonable level of correlation between the predicted and actual values.
Table 3: Result on Continuous Emotion Detection
@@ 6.1 Model performance on continuous emotion detection
6.2 Model comparison on discrete emotion detection
@@
Table 4 shows that the the proposed VAD model stands out with higher precision (0.49) and recall (0.45), suggesting it is more effective in accurately identifying and capturing emotions. ME2E Lite model also performs well, with an F1 score of 0.42, and balanced precision and recall at 0.42 and 0.43, respectively. The baseline ME2E model had the lowest performance, with an F1 score of 0.33 and lower precision and recall (both at 0.32).
@@
Table 4: Result on Discrete Emotion Detection
@@
6.3 Open-vocabulary exploration
At the final stage of our study, we explored the possibility of generating open-vocabulary emotional responses based on six discrete emotion classes. For this, we utilized a subset of 68 samples from the MER2024 dataset which contains the open-vocabulary outputs. We then applied our proposed VAD model to this subset and mapped the resulting VAD labels into the 3D emotion word space showed in the Introduction. To output the final dataset, we used an L2 distance threshold of 0.25, calculated based on the distance which could output an average of five emotion vocabularies from our 3D Emotion Vocabulary Space.
Our finds indicate that our model effectively predicts nuanced emotional states. For instance, in Sample 00000368, the MER2024 dataset listed emotions like ’Alert,’ ’Excited,’ ’Confused,’ and ’Curious,’ and our model predicted ’Shocked.’ This aligns well as ’Shocked’ can encapsulate alertness, excitement, and confusion. In Sample 0002419, the dataset included ’Calm,’ ’Relaxed,’ and ’Happy,’ while our model suggested ’Caring’ and ’Curious.’ These predictions are compatible, with ’Caring’ reflecting a relaxed and content state. Full output list of To further examine the alignment of our open vocabulary, we assess the semantic similarity across various model’s output of emotion labels.
7 Conclusion